# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations in the EU AI Act

### v2

@@ 4. Summary & Outlook: How to proceed?
First, the AI Act heavily focuses on discrimination because of data. The non-discrimination regulations analyzed in the above section mostly focus on bias in (training) data and thus exclude other sources of bias, such as design choices in the algorithms. Furthermore, it is missing specificity on detection and mitigation methods. That makes it hard for practitioners to follow the law.
@@
Second, the AI Act does not include many regulations for none to limited-risk systems. The AI Act not only imposes obligations on high-risk systems but also grants rights, such as the authorization to use special categories of personal data (Article 9 GDPR) in Article 10(5) AIA in high-risk AI contexts. Such permission could also be granted for non-high-risk systems. Furthermore, Such permission could also be granted for non-high-risk systems. Furthermore, non-model specific regulations, such as the obligation to foster AI literacy in Article 4 AIA, are insufficient to ensure non-discriminatory practices. From our perspective, legislators missed an opportunity to regulate AI systems on a broader scope.
@@
Many of the rules focus on (internal) compliance (Zuiderveen Borgesius et al., 2024) and checklist rather than on output-based safeguards for users affected by the AI system. This internal compliance is also linked to the self-assessment approach. It is questionable whether the self-assessment is suitable for protecting individuals and their Fundamental Rights as well as non-discrimination regulation (Bosoer et al., 2023).
@@
Future: A call for more interdisciplinary work. While the AI Act is casuistic and incomplete at the same time (Ződi, 2024), we believe that the AI Act would benefit from four refinements. (1.) If the legislators decide to target non-discrimination regulation, non-discrimination should be explicitly mentioned. In contrast to terms such as bias or fairness, non-discrimination is a well-established legal concept within other European legislative frameworks. (2.) The non-discrimination regulation should cover input as well as output parts of the algorithms in one single article, also along the full AI life cycle (Deck et al., 2024b). Algorithmic output in terms of non-discrimination is only regulated if systems continuously learn after being placed on the market. Otherwise, outputs in terms of discriminatory results of AI systems are not regulated at all. Technical and non-technical means for bias detection, prevention, and mitigation for input and output data should be developed. (3.) Clarification is needed on appropriate measures for non-discrimination. While a reverse burden-of-proof approach is possible here (Wachter, 2024), similar processes as the Code of Practice for GPAI models might be possible such that technical and legal experts as well as all stakeholders work together on refined non-discrimination measures. (4.) If GPAI models are regulated in the current form, a closer link between GPAI models and downstream applications needs to be established (Hacker et al., 2024b). Responsibilities for non-discrimination law compliance must be specified. It is especially important to note that this work is best done in joint community efforts of legal researchers and ML researchers. A bridge between these two fields is highly desirable since regulation should be guided by technical appropriateness and feasibility.
In the introduction of our work, we raised how fair results of AI algorithms can be ensured. We showed that the AI Act tried to solve this non-trivial task with a normative non-discrimination approach. This paper aimed to address legal researchers with an interest in Machine Learning as well as computer scientists and related domains with an interest in legal regulations. We intend to bridge the gap between regulatory and technical limitations in the machine learning domain. From our perspective, both disciplines can benefit from tighter interactions.
5. Acknowledgments

### v3

@@ 4 Summary & Outlook
First, the AI Act focuses on discrimination resulting from data. The non-discrimination regulations analyzed in the above section mostly focus on bias in (training) data and thus exclude other sources of bias, such as design choices in the algorithms. Furthermore, the regulations lack specificity in detection and mitigation methods. That makes it hard for practitioners to follow the law.
@@
Second, the AI Act does not include many regulations for systems ranging from non-risk to limited-risk. The AI Act not only imposes obligations on high-risk systems but also grants rights, such as the authorization to use special categories of personal data (Article 9 GDPR) in Article 10(5) AIA in high-risk AI contexts. Such permission could also be granted for non-high-risk systems. Non-model specific regulations, such as the obligation to foster AI literacy in Article 4 AIA, are insufficient to ensure non-discriminatory practices. From our perspective, legislators missed an opportunity to regulate AI systems on a broader scope.
@@
Many of the rules focus on (internal) compliance (Zuiderveen Borgesius et al. 2024) and checklists rather than on output-based safeguards for users affected by the AI system. This internal compliance is also linked to the self-assessment approach. It is questionable whether the self-assessment is suitable for protecting individuals and their fundamental rights as well as non-discrimination regulation (Bosoer, Cantero Gamito, and Rubio-Marin 2023).
@@
Future. While the AI Act is casuistic and incomplete at the same time (Ződi 2024), we believe that the AI Act would benefit from four refinements. (1.) If the legislators decide to target non-discrimination regulation, non-discrimination should be explicitly mentioned. In contrast to terms such as bias or fairness, non-discrimination is a well-established legal concept within other European legislative frameworks. (2.) For a more consistent interpretation, it would be helpful if non-discrimination regulations cover input as well as output parts of the algorithms in one single article, also along the full AI life cycle (Deck et al. 2024b). Clarification is needed on appropriate measures for non-discrimination. Algorithmic output in terms of non-discrimination is only regulated if systems continuously learn after being placed on the market. Otherwise, outputs in terms of discriminatory results of AI systems are not regulated at all. Technical and non-technical means for bias detection, prevention, and mitigation for input and output data should be developed. For individuals, it could also be important to have a reverse burden-of-proof approach (Wachter 2024), making it easier to claim rights. (3.) Classical EU non-discrimination law will still be important in the AI context. More research is needed on how product safety-related regulations in the AI Act can be complied with by classical individual-based non-discrimination law. (4) Similar to the Code of Practice for GPAI models, it might be possible to start such a process also for high-risk systems. Technical and legal experts, as well as all stakeholders, could work on refined non-discrimination measures for high-risk systems. It is especially important to note that this work is best accomplished through joint community efforts between legal scholars and ML researchers. A bridge between these two fields is highly desirable since regulation should be guided by technical appropriateness and feasibility.
In the introduction of our work, we raised the question of how fair results of AI algorithms can be ensured. While the AI Act provides some direction, we also encounter significant challenges at the intersection of algorithmic fairness and non-discrimination law. Our paper aims to bridge these gaps between regulatory and technical limitations in the machine learning domain. From our perspective, both disciplines can benefit from closer interactions.
5 Acknowledgments